OpenStudy (anonymous):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

Will medal and fan anybody who helps! :) :D <3

OpenStudy (anonymous):

@Squirrels @am!rah @FriedRice @Mathmale

OpenStudy (anonymous):

And everybody leaves .-.

OpenStudy (anonymous):

well do you understand what to do for Part A??

OpenStudy (anonymous):

I dont understand any of this.

OpenStudy (anonymous):

Could you just walk me through it?

OpenStudy (anonymous):

My finals are tommorow, and Im still studying.. xD I just have to know how to do these..

OpenStudy (anonymous):

hmm well you know how Part A is asking you to make up a another equation by replacing one equation with the sum of that equation and a multiple of the other right?

OpenStudy (anonymous):

Yes. I know what they are asking but I don't know how to do it.

OpenStudy (anonymous):

That means that you must solve the equation that is given and then you'll get the answers but you must make a another equation that gives the same answer as the one that you got for the given question

OpenStudy (anonymous):

Can you walk me through it and answer it as we go so I know how to do it on my final?

OpenStudy (anonymous):

This if for part A correct?

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

pretty much

OpenStudy (anonymous):

Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this How would I write what you gave me to fit the answer?

OpenStudy (anonymous):

then now you must make up th equation that gives the answers (0,3/2)

OpenStudy (anonymous):

Can you just answer in Part A: (Answer) and Part B: (Answer)? that way I can submit it on the study guide and get the Extra Credit?

OpenStudy (anonymous):

I closed the question so you can answer :3

OpenStudy (anonymous):

hmmmm but you said to teach you... now you just want me to give you answers?

OpenStudy (anonymous):

Yes.. Please.

OpenStudy (anonymous):

well actually that was the answer to part A

OpenStudy (anonymous):

Can you rewrite it along with part b so I can just copy and paste it and put it in it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Please do this because Ill love you forever.. xD

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

Alrighty! :) <3

OpenStudy (anonymous):

ok this is all combined Part A.)))))))) 3x + 8y = 12 ( 1) 2x + 2y = 3 ( 2) "Replace one equation with the sum of that equation and a multiple of the other." Let us replace (2) by adding (2) to two times (1): 2 times (1): 6x + 16y = 24 Add (2): 2x + 2y = 3 --------------- 8x + 18y = 27 ( 3) So the second pair of equations are: 3x + 8y = 12 ( 1) 8x + 18y = 27 ( 3) Part B>))))))))))) first if you want to find the answer for given question you must first write down 3x + 8y = 12 2x + 2y = 3 on your paper, then you must think if you should use elimination or substitution method well i would use elimination because i do not like fractions... :D 6x + 16y = 24 -6x - 6y = -9 ------------- 0 + 10y = 15 y = 15/10 or 3/2 then you must plug the y into any of the "ORIGINAL" equations.... 2x + 2(3/2) = 3 2x + 3 = 3 2x = 0 x = 0 so the answer for the given question is (0,3/2) 3x + 8y = 12 8x + 18y = 27 -------------- 24x + 64y = 96 -24x - 54y = -81 ---------------- 0 + 10y = 15 y = 3/2 then plug in to "ORIGINAL" equation 3x + 8(3/2) = 12 3x + 12 = 12 3x = 0 x = 0 you will get the answer (0,3/2) as the first given equation

OpenStudy (anonymous):

thats all the work

OpenStudy (anonymous):

and expalination

OpenStudy (anonymous):

so good luck on the exam....

OpenStudy (anonymous):

SO if I copy and paste that I should be good

OpenStudy (anonymous):

you would be more than good its a lot of high quality work

OpenStudy (anonymous):

Do you have to go soon?

OpenStudy (anonymous):

why? you have more?

OpenStudy (anonymous):

I do but one second

OpenStudy (anonymous):

hmmm i wsh its not a long one

OpenStudy (anonymous):

wish*

OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations. graph of line going through ordered pairs (0, 300) and (450, 0)

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

ohh god first make a new chat room

OpenStudy (anonymous):

Wait, then Ill have to close it so you can give me the answers.. lets just stay here?

OpenStudy (anonymous):

hmmmm.... ok

OpenStudy (anonymous):

All I want are answers so I can finish this thing. Dont even bother explaining it.. lol.

OpenStudy (anonymous):

Part A: 2x + 3y = 1,200 y = -2/3x + 400 Part B: by putting the point on the y intercept which is 400 and then use the slope and then draw the line. Part C: f(x) = -2/3x + 400 it represents the profits made by selling phone and tablet cases Part D: you should graph it i cant.... Part E: Well there are parallel functions Part F: i do not know

OpenStudy (anonymous):

thats as far as i can help you

OpenStudy (anonymous):

Is part E answered or no?

OpenStudy (anonymous):

i cant see the graph

OpenStudy (anonymous):

Let me get it for you

OpenStudy (anonymous):

y = -2/3x + 300 well i found out the equation because the y intercept is 300 and the slope was -100/150 = -2/3 which is the slope

OpenStudy (anonymous):

that would be PArt E

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

A couple more questions?!?! I wish I could double medal you! xD

OpenStudy (anonymous):

i have to go sorry

OpenStudy (anonymous):

i wish i could help but i have stuff to do

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Dangit When can you get on next?

OpenStudy (anonymous):

i might be on tommorow

OpenStudy (anonymous):

THank you very much for the help! Youre a life savor!

OpenStudy (anonymous):

around 9 or 10

OpenStudy (anonymous):

np bye

OpenStudy (anonymous):

if you wantyou can fan me so you can know when i join